# Convolutional Neural Networks and Mixture of Experts for Intrusion Detection in 5G Networks and beyond

### v1

@@ VI Conclusion and Future Work
In this paper, we present the first study utilizing Mixture of Experts in the intrusion detection task Specifically, we use a publicly available dataset (5G-NIDD), which is generated from a real 5G test network. Input 1D array of features is transformed into a 2D matrix and then is given as input to CNN layers. The representation vector of the CNN layers is passed through the Mixture of Experts layer consisting of experts and a router. Results showed that our proposed model outperforms state-of-the-art results reaching Accuracy and F1-score up to 0.99958 and 0.99894 respectively. Results of an ablation study demonstrated the effectiveness of all the components of our introduced approach.
Limitations: Our work includes some limitations. Specifically, we used simple methods for imputing missing values, which may influence the evaluation performance. More advanced methods, including Generative Adversarial Imputation Networks (GAIN) [34], have been proposed throughout the years. Additionally, our study relies on labelled dataset.

### v2

@@ VI Conclusion and Future Work
In this paper, we present the first study utilizing Mixture of Experts in the intrusion detection task Specifically, we use a publicly available dataset (5G-NIDD), which is generated from a real 5G test network. Input 1D array of features is transformed into a 2D matrix and then is given as input to CNN layers. The representation vector of the CNN layers is passed through the Mixture of Experts layer consisting of experts and a router. Results on the 5G-NIDD dataset showed that our proposed model outperforms state-of-the-art results reaching Accuracy and F1-score up to 0.99958 and 0.99894 respectively. Results on the NANCY dataset indicate that the proposed approach reaches Accuracy up to 0.78614. Results of an ablation study demonstrated the effectiveness of all the components of our introduced approach.
Limitations: Our work includes some limitations. Specifically, we used simple methods for imputing missing values, which may influence the evaluation performance. More advanced methods, including Generative Adversarial Imputation Networks (GAIN) [35], have been proposed throughout the years. Additionally, our study relies on labelled datasets.